A projectile takes 4secs to pass through points A and B which are at a height 30.4 m above the ground. The Max height attained by the projectile is

Grade:12

3 Answers

Agent K
6 years ago
Let h be max. Height, then Vy (y component of velocity) at h = 0Sy ( distance traveled vertically in those 2 seconds after reaching h max) = 0 + 1/2 g t^2 = 19.6mi.e 19.6 m above 30.4m => H max = 50m
Agent K
6 years ago
Left some spaces... Let h be max. Height, then Vy (y component of velocity) at h = 0Sy ( distance traveled vertically in those 2 seconds after reaching h max) = 0 + 1/2 g t^2 = 19.6mi.e 19.6 m above 30.4m => H max = 50m
Geetha
6 years ago
Max height H =30.4 +x Time taken to pass A and B = 4 secs= time of flightStamderd formula :4H=1/2gt^2 =4x=5*4^2 =x=20m Therefore total height =30+20=50m(Take g=9.8 for more accuracy)